Hi I only need question 6.67.

2 solutions can be found online but the sectioning/cutting is not clear for me.

Through how many joints can we cut through maximum?

A small explanation for every step for and after cutting the truss structure is appreciated thank youCounters Β ΔΟΛ F 6.67 The diagonal members in the center panels of the truss shown are very slender and can act only in tensi
Answer #1

Coun ters BE, CD, DG, EF are Counters u Jokips lo rakips 12kips a & 8or & ft & 8 ft * ett * There is no limit for the numberthe truss, % Let us solve Reqctions: Awy ET v skips K eft x t 8ft 19 kips x 8ft 112 kips oft > EFxzo +) EMA =O → [ Ax =0 → Huour in tension, and ue is positive X - B). For the above FBD) +1 EFy=O 12-6 - FBE Sin O zo 12-6 – FBE Sim (36.8690) =0 | FeeBlow by method of joints, I at joint D, As we know FOX For FDG = Foc to Foc 1 FDG Now +18Fy =o + FDE = 0 FDE 1. Feozo FBE = 1

> Why is CD a zero-force member?

Scuderia Alpha Tauri Honda Fri, Nov 5, 2021 5:51 AM